# Gauging Non-Invertible Symmetries: Topological Interfaces and Generalized Orbifold Groupoid in 2d QFT

### 22 Dec 2023 v2

@@ 2.1 Topological Defect Lines and Fusion Category
¯(Σ)=$\overline{\mathcal{L}}(\Sigma)={\mathcal{L}}(\overline{\Sigma})\,,$ (2.1)
as an operator equation where $\overline{\Sigma}$ has the opposite orientation compared to $\Sigma$. Note that a TDL can be self-dual (such as the identity TDL $\mathds{1}$), i.e. ¯(Σ)=$\overline{\mathcal{L}}(\Sigma)={\mathcal{L}}(\Sigma)$. The TDLs have a natural direct sum structure (denoted by $\oplus$) which amounts to
@@
with fusion coefficients $N_{ij}^{k}\in{\mathbb{Z}}_{\geq 0}$. The fusion product of the simple TDLs then generates a fusion ring. The simple TDLs are ${\mathcal{L}}_{i},{\overline{\mathcal{L}}}_{i}$ such that $N_{i\bar{i}}^{\mathds{1}}=1$
@@
As mentioned in the introduction, a key ingredient of non-invertible symmetry is its intrinsically non-trivial F-symbols that encode a generalized version of the ’t Hooft anomaly for conventional discrete group symmetries. In the context of the above discussion, F-symbols arise from the associativity structure of fusion, which specifies a particular isomorphism, known as the associator ${\alpha}_{i,j,k}$, between the fusion product of three simple TDLs ${\mathcal{L}}_{i},{\mathcal{L}}_{j},{\mathcal{L}}_{j}$ in different orders,
@@
More explicitly, after projecting the triple fusion product to simple TDLs ${\mathcal{L}}_{\ell}$, this is captured by the F-symbol $F_{ijk}^{\ell}$ which can be represented as a matrix (known as the F-matrix) with components $[F_{ijk}^{\ell}]_{(p,{\alpha},{\beta}),(q,{\gamma},{\delta})}$ determining the change of basis between the isomorphic four-point junction vector spaces ${\rm Hom}(({\mathcal{L}}_{i}\otimes{\mathcal{L}}_{j})\otimes{\mathcal{L}}_{k},{\mathcal{L}}_{\ell})$ and ${\rm Hom}({\mathcal{L}}_{i}\otimes({\mathcal{L}}_{j}\otimes{\mathcal{L}}_{k}),{\mathcal{L}}_{\ell})$, where $p,q$ label the simple TDLs ${\mathcal{L}}_{p}\in{\mathcal{L}}_{i}\otimes{\mathcal{L}}_{j},{\mathcal{L}}_{q}\in{\mathcal{L}}_{j}\otimes{\mathcal{L}}_{k}$ in the fusion channels and ${\alpha},{\beta},{\gamma},{\delta}$ label the independent junction vectors at the corresponding trivalent junctions. As usual, these matrix components of the F-symbol are basis dependent in general. Two sets of F-symbols $F_{ijk}^{\ell}$ are equivalent if and only if their matrix components are related by a change of basis at each trivalent junction vector space (i.e. there is a gauge freedom from $[M_{ij}^{k}]_{{\alpha}{\beta}}\in{\rm GL}(N_{ij}^{k},{\mathbb{C}})$ for each triplet of simple TDLs ${\mathcal{L}}_{i},{\mathcal{L}}_{j},{\mathcal{L}}_{k}$).
@@
([${}_{(p,{\alpha},{\beta}),(q,{\gamma},{\delta})}([F_{ijk}^{\ell^{\prime}}]_{(q,{\gamma},{\delta}),(r,\epsilon,\zeta)})^{*}={\delta}_{\ell\ell^{\prime}}{\delta}_{pq}{\delta}_{{\alpha}\epsilon}{\delta}_{{\beta}\zeta}\,.$ (2.10)
@@
The ${\mathcal{C}}$-symmetric CFT enriches the algebraic structure of the symmetry by providing analytical data such as the operator spectrum and correlation functions that transform under the symmetry. In particular, shrinking a TDL ${\mathcal{L}}$ encircling a local operator ${\mathcal{O}}(x)$ produces another local operator ^$\hat{\mathcal{L}}\cdot{\mathcal{O}}(x)$ in the same spacetime representation. The symmetry transformation corresponding to the TDL ${\mathcal{L}}$ is encoded in the linear map $\hat{\mathcal{L}}$ which naturally acts on the CFT Hilbert space ${\mathcal{H}}_{S^{1}}$ on $S^{1}$ by the operator-state correspondence, and it preserves the conformal multiplets labeled by the primary conformal weights $(h,\bar{h})$ with respect to the left and right Virasoro symmetries. The linear maps ^$\hat{\mathcal{L}}_{i}$ for a collection of TDLs ${\mathcal{L}}_{i}$ obviously obey the same fusion rules (on the cylinder). For a CFT with a unique conformally invariant vacuum, which is the focus here, the eigenvalue of ${\mathcal{L}}_{i}$ with respect to the vacuum $|0\rangle$, is the quantum dimension of ${\mathcal{L}}_{i}$,
^$\hat{\mathcal{L}}_{i}|0\rangle=\langle{\mathcal{L}}_{i}\rangle|0\rangle\,.$ (2.11)
@@ 2.2 Algebra Objects and Generalized Gauging
Figure 3: Trivalent topological junction $m$ for the algebra object $A$ which specifies the multiplication morphism and $m^{\dagger}$ for the comultiplication morphism.
@@ 2.3 Half-gauging, Topological Gauging Interfaces and Self-duality
As in the case of discrete group symmetries, instead of gauging an algebra $A\in{\mathcal{C}}$ for a symmetry ${\mathcal{C}}$ of a QFT ${\mathcal{T}}$ on the entire spacetime, we can consider gauging on a submanifold ${\mathcal{S}}$ of the spacetime. In particular, by choosing a topological boundary condition for the relevant gauge fields at the boundary $\partial{\mathcal{S}}$, this procedure produces a topological interface ${\cal I}_{T|{\mathcal{T}}/A}({\mathcal{S}})$ (which can be disconnected if ${\mathcal{S}}$ is) located at ${\mathcal{S}}$ interpolating between the original theory ${\mathcal{T}}$ and its gauged cousin ${\mathcal{T}}/A$. Due to the topological nature of this interface, to study its properties it suffices to focus on the simplest setup where ${\mathcal{S}}$ is half of the spacetime as in Figure 5.
@@
Figure 6: Fusion of topological gauging interfaces from half-gauging $A$.
@@
producing a (non-simple) TDL in the theory ${\mathcal{T}}/A$, which we denote as $A^{*}$, that must have the same quantum dimension as $A$
@@
${\mathcal{N}}\otimes\overline{{\mathcal{N}}}=A\,,\quad\overline{{\mathcal{N}}}\otimes{{\mathcal{N}}}=A^{*}\,,\quad$ (2.20)
where the isomorphism in (2.18) is implicitly used to identify TDLs in ${\mathcal{T}}/A$ with those in ${\mathcal{T}}$ [^10] The full set of fusion rules involving the simple TDLs in $A$ and $A^{*}$ with ${\mathcal{N}}$ and $\overline{{\mathcal{N}}}$ will be captured by the category of right and left $A$-modules, respectively (physically these correspond to other possible simple topological interfaces between ${\mathcal{T}}$ and ${\mathcal{T}}/A$). See Section 2.4 for related discussions. Note that, in general $A$ and $A^{*}$ are non-isomorphic as TDLs. Furthermore, the F-symbols involving the TDL ${\mathcal{N}}$ (and its dual) will depend on the multiplication morphism $m$ for the corresponding algebra $A$ that implements the self-dual gauging.
@@ 2.4 General Topological Interfaces and (Bi)module Categories
where $P_{ia}{}^{b}$ has non-negative integer entries and counts the number of independent topological junctions between ${\mathcal{L}}_{i},{\cal I}_{a}$ and ${\cal I}_{b}$. There is an analog of the F-matrix for the change of basis matrix in junction vector spaces between bulk TDLs and interfaces. Consistency from isotopy invariance in the presence of multiple topological junctions along an interface leads to an interface analog of the pentagon equation (see [56] for a recent review). The full mathematical structure (up to equivalence from basis change at the junction vector spaces) is captured by a left ${\mathcal{C}}$-module category, which we denote as ${\mathcal{M}}\in{\rm Mod}({\mathcal{C}})$. In this language, the simple interfaces ${\cal I}_{a}$ are the simple objects in ${\mathcal{M}}$, the topological junction vector space between two general interfaces ${\cal I},{\cal I}^{\prime}$ (from direct sums of ${\cal I}_{a}$) is identified with the Hom space ${\rm Hom}_{\mathcal{M}}({\cal I},{\cal I}^{\prime})$. In particular, the NIM-rep gives
@@
More generally, a $({\mathcal{C}},{\mathcal{C}}^{\prime})$-bimodule category for fusion categories ${\mathcal{C}}$ and ${\mathcal{C}}^{\prime}$ represents a category of topological interfaces where TDLs in ${\mathcal{C}}$ and ${\mathcal{C}}^{\prime}$ can end topologically from the left and from the right respectively. The bimodule categories have a natural tensor product from the fusion of the topological interfaces, which is mathematically defined as ${\mathcal{M}}\boxtimes_{{\mathcal{C}}^{\prime}}{\mathcal{N}}$ for a $({\mathcal{C}},{\mathcal{C}}^{\prime})$-bimodule category ${\mathcal{M}}$ and a $({\mathcal{C}}^{\prime},{\mathcal{C}}^{\prime\prime})$-bimodule category ${\mathcal{N}}$ [71]. A $({\mathcal{C}},{\mathcal{C}}^{\prime})$-bimodule category ${\mathcal{M}}$ is called invertible if ${\mathcal{M}}^{\rm op}\boxtimes_{{\mathcal{C}}}{\mathcal{M}}\cong{\mathcal{C}}^{\prime}$ (equivalently ${\mathcal{M}}\boxtimes_{{\mathcal{C}}^{\prime}}{\mathcal{M}}^{\rm op}\cong{\mathcal{C}}$) as bimodule categories where ${\mathcal{M}}^{\rm op}$ is the $({\mathcal{C}}^{\prime},{\mathcal{C}})$-bimodule category defined by orientation reversal. Physically, any simple object in an invertible bimodule is a topological interface ${\cal I}$ such that the fusion product contains the trivial TDL without degeneracy ¯${\cal I}\otimes\overline{{\cal I}}\ni\mathds{1}_{\mathcal{C}}$ and ¯$\overline{{\cal I}}\otimes{\cal I}\ni\mathds{1}_{{\mathcal{C}}^{\prime}}$. When this happens, the two fusion categories ${\mathcal{C}}$ and ${\mathcal{C}}^{\prime}$ are said to be categorically Morita equivalent [71]. The $({\mathcal{C}},{}_{A}{\mathcal{C}}_{A})$-bimodule category ${\mathcal{M}}={\mathcal{C}}_{A}$ obtained from gauging is precisely of this type, where ${\cal I}$ is the topological gauging interface for the haploid algebra $A={\cal I}\otimes\overline{{\cal I}}$ and $A^{*}=\overline{{\cal I}}\otimes{\cal I}$ is the dual algebra object. The $A$-gauging can be undone by further gauging $A^{*}$. In fact, two fusion categories ${\mathcal{C}}$ and ${\mathcal{C}}^{\prime}$ are categorically Morita equivalent if and only if each is the dual of the other by gauging.
The categorical Morita equivalence is in fact a 2-equivalence between certain 2-categories [69, 71, 95, 11]. Here the 2-category ${\rm Mod}({\mathcal{C}})$ contains as objects module categories ${\mathcal{M}},{\mathcal{N}}$ over a fusion category ${\mathcal{C}}$, the 1-morphisms are given by ${\mathcal{C}}$-module functors ${\rm Fun}_{\mathcal{C}}({\mathcal{M}},{\mathcal{N}})$, and the 2-morphisms by the natural transformations of ${\mathcal{C}}$-module functors. Equivalently, the simple objects in ${\rm Mod}({\mathcal{C}})$ can be thought of as the set of haploid algebras $A\in{\mathcal{C}}$ (referred to as division algebras in [95]) up to Morita equivalence, while the 1-morphisms and the 2-morphisms are given by objects and morphisms respectively in the category of $(A,B)$-bimodules. In physical terms, the objects ${\mathcal{M}},{\mathcal{N}}$ label distinct gaugings of TDLs in the symmetry subcategory ${\mathcal{C}}$ of the theory ${\mathcal{T}}$, the 1-morphisms correspond to topological interfaces between (potentially) different gauged theories, and the 2-morphisms describe topological interface changing operators. This 2-equivalence implies a bijection ${\mathcal{N}}\to{\rm Fun}_{\mathcal{C}}({\mathcal{M}},{\mathcal{N}})$ between the module categories over a pair of fusion categories ${\mathcal{C}}$ and ${}_{A}{\mathcal{C}}_{A}$ that are dual with respect to the ${\mathcal{C}}$-module category ${\mathcal{M}}$. In particular, for ${\mathcal{N}}={\mathcal{M}}$, this implies the isomorphism ${}_{A}{\mathcal{C}}_{A}\cong({\mathcal{C}}^{*}_{\mathcal{M}})^{\rm op}$ as fusion categories with ${\mathcal{C}}^{*}_{\mathcal{M}}\equiv{\rm Fun}_{\mathcal{C}}({\mathcal{M}},{\mathcal{M}})$.
@@
The fusion categories related by discrete gauging in this way form a graph where each connected component consists of nodes labeled by fusion categories that are categorically Morita equivalent and connecting edges corresponding to inequivalent invertible bimodule categories. This is the generalized version the orbifold groupoid introduced in [46] where it was studied in detail for invertible gaugings. [^16] A slight difference with [46] is that self-dual gaugings do not create a separate node in the definition here (and in [11]). This generalized orbifold groupoid structure is mathematically captured by the Brauer-Picard groupoid $\underline{\underline{\rm\bf BrPic}}$; a 3-groupoid which is a special case of a 3-category whose objects are fusion categories, 1-morphisms are invertible $({\mathcal{C}},{\mathcal{C}}^{\prime})$-bimodules between fusion categories ${\mathcal{C}}$ and ${\mathcal{C}}^{\prime}$, 2-morphisms are equivalences between such invertible bimodules, and 3-morphisms are isomorphisms of such equivalences (see [11] for more details). Here we focus on the truncated 1-groupoid which is a connected subgroupoid that contains a seed fusion category ${\mathcal{C}}$ which is denoted simply as ${\rm\bf BrPic}$ (where the ${\mathcal{C}}$ dependence is implicit [11]). Obviously, the definition of this groupoid is independent of the fusion category up to categorical Morita equivalence. Physically, the groupoid ${\rm\bf BrPic}$ keeps track of the fusion structure of topological interfaces described by invertible bimodule categories over fusion categories that are related by discrete gauging (generalized orbifold) to ${\mathcal{C}}$. As we will illustrate in examples in Section 4, this immediately predicts the fate of sequential gauging in a ${\mathcal{C}}$-symmetry theory ${\mathcal{T}}$ without knowing details of the dynamical content of the theory (i.e. it tells when seemingly different gauging sequences are equivalent). In particular, the Morita auto-equivalences of ${\mathcal{C}}$ (i.e. invertible $({\mathcal{C}},{\mathcal{C}})$-bimodule categories) form a group denoted by ${\rm\bf BrPic}({\mathcal{C}})$ and captures potential nontrivial isomorphisms of ${\mathcal{C}}$ [^17] We emphasize that while the underlying QFT ${\mathcal{T}}$ retains the ${\mathcal{C}}$-symmetry after the discrete gauging which corresponds to the $({\mathcal{C}},{\mathcal{C}})$-bimodule category, in general ${\mathcal{T}}$ is not necessarily self-dual under this gauging. Such nontrivial isomorphism can come from an automorphism of the fusion category (e.g. a permutation of TDLs that respect the fusion rules) or stacking with a topological counter-term, equivalently an SPT for ${\mathcal{C}}$.

### 25 Mar 2026 v3

@@ 2.1 Topological Defect Lines and Fusion Category
$\overline{{\mathcal{L}}}(\Sigma)={\mathcal{L}}(\overline{\Sigma})\,$, (2.1)
as an operator equation where $\overline{\Sigma}$ has the opposite orientation compared to $\Sigma$. Note that a TDL can be self-dual (such as the identity TDL $\mathds{1}$), i.e. $\overline{{\mathcal{L}}}(\Sigma)={\mathcal{L}}(\Sigma)$. The TDLs have a natural direct sum structure (denoted by $\oplus$) which amounts to
@@
with fusion coefficients $N_{ij}^{k}\in{\mathbb{Z}}_{\geq 0}$. The fusion product of the simple TDLs then generates a fusion ring. The simple TDLs are ${\mathcal{L}}_{i},{\overline{{\mathcal{L}}}}_{i}$ such that $N_{i\bar{i}}^{\mathds{1}}=1$ .
@@
As mentioned in the introduction, a key ingredient of non-invertible symmetry is its intrinsically non-trivial F-symbols that encode a generalized version of the ’t Hooft anomaly for conventional discrete group symmetries. In the context of the above discussion, F-symbols arise from the associativity structure of fusion, which specifies a particular isomorphism, known as the associator ${\alpha}_{i,j,k}$, between the fusion product of three simple TDLs ${\mathcal{L}}_{i},{\mathcal{L}}_{j},{\mathcal{L}}_{k}$ in different orders,
@@
More explicitly, after projecting the triple fusion product to simple TDLs $\mathcal{L}$$_{\ell}$, this is captured by the F-symbol $F_{ijk}^{\ell}$ which can be represented as a matrix (known as the F-matrix) with components $F_{ijk}^{\ell}]_{(p,{\alpha},{\beta}),(q,{\gamma},{\delta})}$ determining the change of basis between the isomorphic four-point junction vector spaces ${\rm Hom}(({\mathcal{L}}_{i}\otimes{\mathcal{L}}_{j})\otimes{\mathcal{L}}_{k},{\mathcal{L}}_{\ell})$ and ${\rm Hom}({\mathcal{L}}_{i}\otimes({\mathcal{L}}_{j}\otimes{\mathcal{L}}_{k}),{\mathcal{L}}_{\ell})$, where $p,q$ label the simple TDLs ${\mathcal{L}}_{p}\in{\mathcal{L}}_{i}\otimes{\mathcal{L}}_{j},{\mathcal{L}}_{q}\in{\mathcal{L}}_{j}\otimes{\mathcal{L}}_{k}$ in the fusion channels and ${\alpha},{\beta},{\gamma},{\delta}$ label the independent junction vectors at the corresponding trivalent junctions. As usual, these matrix components of the F-symbol are basis dependent in general. Two sets of F-symbols $F_{ijk}^{\ell}$ are equivalent if and only if their matrix components are related by a change of basis at each trivalent junction vector space (i.e. there is a gauge freedom from $M_{ij}^{k}]_{{\alpha}{\beta}}\in{\rm GL}(N_{ij}^{k},{\mathbb{C}})$ for each triplet of simple TDLs ${\mathcal{L}}_{i},{\mathcal{L}}_{j},{\mathcal{L}}_{k}$).
@@
${}_{(p,{\alpha},{\beta}),(q,{\gamma},{\delta})}([F_{ijk}^{\ell^{\prime}}]_{(q,{\gamma},{\delta}),(r,\epsilon,\zeta)})^{*}={\delta}_{\ell\ell^{\prime}}{\delta}_{pq}{\delta}_{{\alpha}\epsilon}{\delta}_{{\beta}\zeta}\,.$ (2.10)
@@
The $\mathcal{C}$}-symmetric CFT enriches the algebraic structure of the symmetry by providing analytical data such as the operator spectrum and correlation functions that transform under the symmetry. In particular, shrinking a TDL $\mathcal{L}$} encircling a local operator ${\mathcal{O}}(x)$ produces another local operator $\hat{\mathcal{L}}\cdot{\mathcal{O}}(x)$ in the same spacetime representation. The symmetry transformation corresponding to the TDL $\mathcal{L}$} is encoded in the linear map $\hat{\mathcal{L}}$ which naturally acts on the CFT Hilbert space ${\mathcal{H}}_{S^{1}}$ on $S^{1}$ by the operator-state correspondence, and it preserves the conformal multiplets labeled by the primary conformal weights $(h,\bar{h})$ with respect to the left and right Virasoro symmetries. The linear maps $\hat{\mathcal{L}}_{i}$ for a collection of TDLs ${\mathcal{L}}_{i}$ obviously obey the same fusion rules (on the cylinder). For a CFT with a unique conformally invariant vacuum, which is the focus here, the eigenvalue of ${\mathcal{L}}_{i}$ with respect to the vacuum $|0\rangle$, is the quantum dimension of ${\mathcal{L}}_{i}$,
$\hat{\mathcal{L}}_{i}|0\rangle=\langle{\mathcal{L}}_{i}\rangle|0\rangle\,.$ (2.11)
@@ 2.2 Algebra Objects and Generalized Gauging
Figure 3: Trivalent topological junction $m$ for the algebra object AA which specifies the multiplication morphism and $m^{\dagger}$ for the comultiplication morphism.
@@ 2.3 Half-gauging, Topological Gauging Interfaces and Self-duality
As in the case of discrete group symmetries, instead of gauging an algebra $A\in{\mathcal{C}}$ for a symmetry $\mathcal{C}$} of a QFT $\mathcal{T}$} on the entire spacetime, we can consider gauging on a submanifold $\mathcal{S}$} of the spacetime. In particular, by choosing a topological boundary condition for the relevant gauge fields at the boundary $\partial{\mathcal{S}}$, this procedure produces a topological interface ${\cal I}_{{\mathcal{T}}|{\mathcal{T}}/A}({\mathcal{S}})$ (which can be disconnected if $\mathcal{S}$} is) located at $\mathcal{S}$} interpolating between the original theory $\mathcal{T}$} and its gauged cousin ${\mathcal{T}}/A$. Due to the topological nature of this interface, to study its properties it suffices to focus on the simplest setup where $\mathcal{S}$} is half of the spacetime as in Figure 5.
@@
Figure 6: Fusion of topological gauging interfaces from half-gauging AA.
@@
producing a (non-simple) TDL in the theory ${\mathcal{T}}/A$, which we denote as $A^{*}$, that must have the same quantum dimension as AA
@@
${\mathcal{N}}\otimes\overline{{\mathcal{N}}}=A\,,\quad\overline{{\mathcal{N}}}\otimes{{\mathcal{N}}}=A^{*}\,$$\quad$ (2.20)
where the isomorphism in (2.18) is implicitly used to identify TDLs in ${\mathcal{T}}/A$ with those in $\mathcal{T}$}. [^10] The full set of fusion rules involving the simple TDLs in $A$ and $A^{*}$ with $\mathcal{N}$} and $\overline{{\mathcal{N}}}$ will be captured by the category of right and left $A$-modules, respectively (physically these correspond to other possible simple topological interfaces between $\mathcal{T}$} and ${\mathcal{T}}/A$). See Section 2.4 for related discussions. Note that, in general $A$ and $A^{*}$ are non-isomorphic as TDLs. Furthermore, the F-symbols involving the TDL $\mathcal{N}$} (and its dual) will depend on the multiplication morphism $m$ for the corresponding algebra $A$ that implements the self-dual gauging.
@@ 2.4 General Topological Interfaces and (Bi)module Categories
[…]
@@
[…]
The categorical Morita equivalence is in fact a 2-equivalence between certain 2-categories [69, 71, 95, 11]. Here the 2-category ${\rm Mod}({\mathcal{C}})$ contains as objects module categories ${\mathcal{M}},{\mathcal{N}}$ over a fusion category $\mathcal{C}$}, the 1-morphisms are given by $\mathcal{C}$}-module functors ${\rm Fun}_{\mathcal{C}}({\mathcal{M}},{\mathcal{N}})$, and the 2-morphisms by the natural transformations of $\mathcal{C}$}-module functors. Equivalently, the simple objects in ${\rm Mod}({\mathcal{C}})$ can be thought of as the set of haploid algebras $A\in{\mathcal{C}}$ (referred to as division algebras in [95]) up to Morita equivalence, while the 1-morphisms and the 2-morphisms are given by objects and morphisms respectively in the category of $(A,B)$-bimodules. In physical terms, the objects ${\mathcal{M}},{\mathcal{N}}$ label distinct gaugings of TDLs in the symmetry subcategory $\mathcal{C}$} of the theory $\mathcal{T}$}, the 1-morphisms correspond to topological interfaces between (potentially) different gauged theories, and the 2-morphisms describe topological interface changing operators. This 2-equivalence implies a bijection ${\mathcal{N}}\to{\rm Fun}_{\mathcal{C}}({\mathcal{M}},{\mathcal{N}})$ between the module categories over a pair of fusion categories $\mathcal{C}$} and $_{A}{\mathcal{C}}_{A}$ that are dual with respect to the $\mathcal{C}$}-module category $\mathcal{M}$}. In particular, for $\mathcal{N}$$\mathcal{M}$}, this implies the isomorphism $_{A}{\mathcal{C}}_{A}\cong({\mathcal{C}}^{*}_{\mathcal{M}})^{\rm op}$ as fusion categories with ${\mathcal{C}}^{*}_{\mathcal{M}}\equiv{\rm Fun}_{\mathcal{C}}({\mathcal{M}},{\mathcal{M}})$.
@@
The fusion categories related by discrete gauging in this way form a graph where each connected component consists of nodes labeled by fusion categories that are categorically Morita equivalent and connecting edges corresponding to inequivalent invertible bimodule categories. This is the generalized version of the orbifold groupoid introduced in [46] where it was studied in detail for invertible gaugings. [^16] A slight difference with [46] is that self-dual gaugings do not create a separate node in the definition here (and in [11]). This generalized orbifold groupoid structure is mathematically captured by the Brauer-Picard groupoid $\underline{\underline{\rm\bf BrPic}}$; a 3-groupoid which is a special case of a 3-category whose objects are fusion categories, 1-morphisms are invertible $({\mathcal{C}},{\mathcal{C}}^{\prime})$-bimodules between fusion categories $\mathcal{C}$} and $\mathcal{C}$$^{\prime}$, 2-morphisms are equivalences between such invertible bimodules, and 3-morphisms are isomorphisms of such equivalences (see [11] for more details). Here we focus on the truncated 1-groupoid which is a connected subgroupoid that contains a seed fusion category $\mathcal{C}$} which is denoted simply as $\rm\bf BrPic$} (where the $\mathcal{C}$} dependence is implicit [11]). Obviously, the definition of this groupoid is independent of the fusion category up to categorical Morita equivalence. Physically, the groupoid $\rm\bf BrPic$} keeps track of the fusion structure of topological interfaces described by invertible bimodule categories over fusion categories that are related by discrete gauging (generalized orbifold) to $\mathcal{C}$}. As we will illustrate in examples in Section 4, this immediately predicts the fate of sequential gauging in a $\mathcal{C}$}-symmetry theory $\mathcal{T}$} without knowing details of the dynamical content of the theory (i.e. it tells when seemingly different gauging sequences are equivalent). In particular, the Morita auto-equivalences of $\mathcal{C}$} (i.e. invertible $({\mathcal{C}},{\mathcal{C}})$-bimodule categories) form a group denoted by ${\rm\bf BrPic}({\mathcal{C}})$ and captures potential nontrivial isomorphisms of $\mathcal{C}$}. [^17] We emphasize that while the underlying QFT $\mathcal{T}$} retains the $\mathcal{C}$}-symmetry after the discrete gauging which corresponds to the $({\mathcal{C}},{\mathcal{C}})$-bimodule category, in general $\mathcal{T}$} is not necessarily self-dual under this gauging. Such nontrivial isomorphism can come from an automorphism of the fusion category (e.g. a permutation of TDLs that respect the fusion rules) or stacking with a topological counter-term, equivalently an SPT for $\mathcal{C}$}.